OpenStudy (anonymous):

Trig identity question..... Write the first expression in terms of the second if the terminal point determined by t is in the given quadrant. sin t, sec t; Quadrant IV

OpenStudy (valpey):

In quadrant IV, sin is negative and cosine is positive. secant is the reciprocal of cosine so is also positive. The pertinent Pythagorean identity is \[\sin^2{t} = 1-\cos^2{t}\]Which is equal to \[1-\frac{1}{\sec^2{t}}\] We can take the square root to get sin(t), but we should add the negative sign because of the quadrant: \[\sin{t}=-\sqrt{1-\frac{1}{\sec^2{t}}}\]

OpenStudy (anonymous):

thank you for explaining and not just an answer.... may have two more...... just going to try and work them out myself first.

OpenStudy (valpey):

Good luck

OpenStudy (valpey):

I should mention that we would have the same result in Quadrant III because it doesn't really matter what the sign of sec(t) is, just the sin(t).

OpenStudy (anonymous):

here's another.. Write the first expression in terms of the second if the terminal point determined by t is in the given quadrant. tan t, cos t; Quadrant III I got \[\tan ^{2}t=1-\sin ^{2}t/\cos ^{2}t\] but says its wrong. am i just missing a negative sign?